THE SUM OF A NUMBER AND TWICE ITS RECIPROCAL IS 9/2. fIND THE NUMBER
Answer by checkley75(3666)

X+2*1/X=9/2
X+2/X=9/2
(X^2+2)/X=9/2 NOW CROSS MULTIPLY
2(X^2+2)=9X
2X^2+4-9X=0
2X62-9X+4=0
(2X-1)(X-4)=0
2X-1=0
2X=1
X=1/2 ANSWER.
X-4=0
X=4 ANSWER.